OpenStudy (anonymous):

Soybean meal is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb miixture that is 11% protein? How do I even start this?

OpenStudy (anonymous):

soybean + Cornmeal = 360 lb .18soybean + .09Cornmeal = .11(360)

OpenStudy (anonymous):

and then soybean = 360 - Cornmeal , plug this into the bottom equation

OpenStudy (callisto):

let x pounds be the pounds of Soybean meal , then, pounds of cornmeal = 360- x 18% (x) + (360-x) 9% = 11% (360) 0.18x-0.09x + 32.4 = 39.6 x=80 therefore pounds of Soybean meal required is 80 pounds and pounds of cornmeal required is (360-80)pounds, that is 280 pounds

OpenStudy (anonymous):

.18(360-Cornmeal) + .09Cornmeal = .11(360)

OpenStudy (anonymous):

Nevermind Callisto is Faster =)

OpenStudy (callisto):

i thought i was the slower one XD